C96C2

PI: Rioja/Sanghera
Title: CAL - DA193

Session: June 1996 (session 2)
Frequency: 6 cm
Mode: A (dual polarisation)
Obs time (UT): 167 EU 08:00(15/06)-09:00(15/06)
Obs time (UT): 170 EU 08:00(18/06)-09:00(18/06)
Stations: JbNtOnWbToSmUrSh
Processor: Bonn
EU Mk3 Tapes/St.: 1

Comment


To give feedback on this experiment, click here

Experiment Feedback

Onsala (John Conway at Sun Jun 16 03:33:02 EDT 1996) Success
Part 1. No known problesm

Torun (Roman Feiler at Sun Jun 16 08:03:57 EDT 1996) Success
Only LHC polarisation recorded - 8 BBC

Noto (Corrado Trigilio at Tue Jun 18 09:18:27 EDT 1996) Minor failures
NEWS in Noto: phase-cal for the 6cm receiver; GPS receiver available. 12 BBCs avaliable (still missing bbc12 and bbc14) Part 1) No problems reported; good sky conditions Part 2) first scan (pass 5) partially lost for Field-System crash due to mat communications problems during Tsys measurement (3 min lost). No other problems reported. Good sky conditions.

Onsala (Johm Conway at Tue Jun 18 09:23:38 EDT 1996) Success
Part 2 - high winds (12m/s) degraded pointing somewhat, otherwise no known problesm.

Westerbork (tony foley at Thu Jun 20 05:46:37 EDT 1996) Minor failures
Day 167 Used the tied-array with 3 telescopes (RTs 8,9,A) and phasecal injection into tied-array (via input for RT0) Day 170 Before the observation the formatter day number had to be corrected as it did not increment at 0UT. We used a single-dish setup (RT8) and phasecal injection coupled into the X dipole of RT8. Telescope pointing via the Field System..

Westerbork (foley at Thu Jun 20 05:50:42 EDT 1996) Minor failures
ALSO problem with 50millihertz offset in the tied-array frequency on day 167

Jodrell (David Henstock at Fri Jun 21 08:54:48 EDT 1996) Success
No problems with CAL experiment.

Urumqi (Lee Suqin at Tue Jul 2 01:52:08 EDT 1996) Success
Success on day 170; It isclear. Clock offset =0.6 /-0.5us.

Urumqi (Lee Suqin at Tue Jul 2 02:11:56 EDT 1996) Failure
On day 167 It was rain; The comunication between the tape-recorder and the computer was not good. the data maybe incorrect. The clock offset =0.2 /- 0.5us. Not set IFD=46 during 8:00-8:13 scan. No phase cal for 6cm receiver. Tsys=313K; effective factor=49.5%; cal temp =67K.

Shanghai (Xiao\-Yu Tan at Mon Jul 15 12:41:30 EDT 1996) Success
Date of tapes shipment was 21, June 1996. All obs took place. No problems to report. Only LCP data is available. Clock offset on each of the two days was -0.6 and -0.5 micro sec on days 167 and 170 respectively.